# Emergence of Gibbs ensemble as the unique steady state in open quantum systems

### v1

@@ Abstract
We explicitly construct unique non-equilibrium steady state (NESS) of Lindblad master equation characterized by a Gibbs ensemble $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$, where the effective Hamiltonian $\tilde{H}$ consists only of $U(1)$ conserved charges of the original Hamiltonian. Specifically, when the original Hamiltonian has multiple charges, it is possible to couple them with bathes at different temperature respectively, but still leads to an equilibrium state. To access the Gibbs NESS, the jump operators need to be properly chosen to fulfill quantum detailed balance condition (qDBC). These jump operators are ladder operators for $\tilde{H}$ and jump process they generate form a vertex-weighted directed acyclic graph (wDAG). By studying the XX model and Fredkin model, we showcase how the Gibbs state emerges as the unique steady state.
Introduction.— The Lindblad master equation is a paradigm for describing Markovian open quantum systems (Gorini et al., 1976; Lindblad, 1976; Breuer and Petruccione, 2007). It represents the dynamics as an average of different trajectories, which consist of non-Hermitian evolution and stochastic quantum jumps, also known as "unraveling." In master equation form, the dynamics generator is
@@
where $L_{k}$s are jump operators and $H_{\text{eff}}=H-\frac{i}{2}\sum_{k}L_{k}^{\dagger}L_{k}$. The detailed balance condition (DBC) is not required in Lindblad dynamics, making it powerful and capable of characterizing non-equilibrium phenomena (Kamenev, 2011; Derrida, 2007; Garrahan, 2018; Prosen, 2008; Verstraete et al., 2009; Dubois et al., 2023; Liu et al., 2024; Buča et al., 2019; Rakovszky et al., 2023; Wang et al., 2023; Haga et al., 2021; Schmolke and Lutz, 2022; Zhou et al., 2022; Zheng et al., 2023). Recently with development of quantum computation, people regain interests in crafting quantum Gibbs samplers and preparing specific state utilizing Lindblad dynamics (Rall et al., 2023; Chen et al., 2023; Ding et al., 2024; Guo et al., 2024). In this work we make direct use of qDBC to derive exact Gibbs state that only depends on $U(1)$ conserved charges of original Hamiltonian.
DBC originates from classical statistical physics and guarantees an equilibrium stationary state. In a classical Markov process generated by W, the probability distribution of configuration set {c} satisfies $\frac{\mathrm{d}P(c)}{\mathrm{d}t}=\sum_{c^{\prime}\neq c}W(c^{\prime}\rightarrow c)P(c^{\prime})-W(c\rightarrow c^{\prime})P(c)$. For a stationary state satisfying classical DBC, $W(c^{\prime}\rightarrow c)P(c^{\prime})=W(c\rightarrow c^{\prime})P(c)$. DBC can also be interpreted as time-reversal symmetry, as it constrains the relative probabilities of configurations connected by a Markov process. In other words, starting from an arbitrary configuration, one can determine the probability of any other configuration without frustration. Besides the above intuitive condition, Kubo-Martin-Schwinger (KMS) condition (Haag et al., 1967) $\langle\hat{A}(t)\hat{B}(0)\rangle_{\text{eq}}=\langle\hat{B}(t+i\beta)\hat{A}(0)\rangle_{\text{eq}}$ also acts as definition for thermal equilibrium state. It suggests a generalized "time evolution" acting on $\hat{B}$. qDBC extends this idea by requiring that this "time evolution" is also a completely positive and trace-preserving (CPTP) map, or a quantum Markov semigroup (QMS). Similar idea is applied to study properties of thermal states (Firanko et al., 2024) and mixing time (Bardet et al., 2023), to solve some models exactly (de Leeuw et al., 2024; Roberts et al., 2021; Roberts and Clerk, 2023) and a local thermal state is constructed (Palmero et al., 2019). In the context of Lindblad dynamics $\mathcal{L}(H,\{L_{k}\})$, the qDBC condition for an invertible (faithful) density matrix $\rho$ manifests as (Alicki, 1976; Fagnola and Umanità, 2007)
@@
where $\tilde{L}_{k}=UL_{k}$, $U$ is an unitary that operates on space spanned by {$L_{k}$}. Liouvillian of this form that hosts a Gibbs NESS is often denoted as Davies generator (Davies, 1974).
@@
When jump operators are Hermitian, the identity matrix is always a NESS due to hermiticity of the Liouvillian. However, when jump operators are non-Hermitian and the Liouvillian is not diagonal-preserving, the NESS is typically not a diagonal thermal equilibrium state and usually has finite stationary current, especially in the case of boundary driving (Prosen, 2015). Nevertheless, in this work we demonstrate that a diagonal thermal equilibrium state can be constructed as a NESS in various models. Such a state satisfies qDBC, making it a thermal equilibrium state that depends only on the conserved quantities in the original Hamiltonian.
@@
Our method introduced here relies on the coherent Hamiltonian having conservation laws, such that its Hilbert space is fractured into some disconnected pieces, which we denote as Krylov sectors or dynamical sectors. Besides conventional $U(1)$ conversational laws like particle conservation (Buča and Prosen, 2012), it is also applicable to non-Abelian cases (Zhang et al., 2020), dynamical constrained systems (Ritort and Sollich, 2003; Fredrickson and Andersen, 1984) and systems with Hilbert space fragmentation (HSF) (Sala et al., 2020; Moudgalya and Motrunich, 2022; Li et al., 2023; Balasubramanian et al., 2024).
@@
General theory.— Throughout this work we only consider separable Hilbert space with finite local dimension, such as spin system or fermion system. Our goal is to achieve a hierarchical structure of NESS that resembles equilibrium Gibbs distribution. Since a density matrix $\rho$ is positive and Hermitian, it can be expressed as $\rho=e^{-\beta\tilde{H}}$, where $\tilde{H}$ is a Hermitian matrix and $\beta\in\mathbb{R}$. To clarify Gibbs state, we construct $\tilde{H}$ by conserved quantities in original Hamiltonian and relate $\beta$ to temperature-alike quantity as in classical statistical physics. To realize this hierarchy, we pack states by taking conserved quantities as our packing principle. Since conserved quantities separate Hilbert space into disconnected sectors, every state sharing the same quantum numbers should have the same weight in $\rho$. An effective tool to characterize these conservation laws is commutant algebra. The commutant algebra formalism, as discussed in (Moudgalya and Motrunich, 2022), is particularly useful for analyzing HSF, but in general it is not restricted to HSF cases. Commutant algebra formalism regards every invariant subspace of Hamiltonian as representation space for each irreducible representation (irrep) of bond algebra $\mathcal{A}$ generated by Hamiltonian terms $\{h_{\text{loc}}\}$ and identity operator $I$. The commutant algebra $\mathcal{C}$ consists of all operators $\{O\}$ of Hilbert space $\mathcal{H}$ that commute with all generators of $\mathcal{A}$, i.e., $[h_{\text{loc},i},O]=0$ for all $i$. The total Hilbert space can be decomposed by irreps of bond algebra and commutant algebra marked by $\lambda$, i.e., $\mathcal{H}=\bigoplus_{\lambda}\mathcal{H}^{\mathcal{A}}_{\lambda}\otimes\mathcal{H}^{\mathcal{C}}_{\lambda}$. If $\mathcal{H}^{\mathcal{C}}_{\lambda}$ has dimension $d$ larger than 1, then the Hamiltonian has degenerated sectors $H_{\lambda,1,...d}$. An operator in $\mathcal{A}$ has form $\oplus_{\lambda}h_{\lambda}\otimes\mathcal{I}$, and an operator in $\mathcal{C}$ has form $\oplus_{\lambda}\mathcal{I}\otimes c_{\lambda}$ where $h_{\lambda}$ and $c_{\lambda}$ is supported on $\mathcal{H}^{\mathcal{A}}_{\lambda}$ and $\mathcal{H}^{\mathcal{C}}_{\lambda}$, respectively.
@@
Now returning to qDBC (2), natural questions arise: how is $\tilde{H}$ related to conserved quantities of the original Hamiltonian and how can we find specific jump operators that satisfies qDBC? We will see they are related. First let us consider choice of jump operators. If $\lambda_{k}=1$, then a direct choice for $\tilde{L}_{k}$ to satisfy qDBC will be an element of bond algebra or a diagonal operator like $S^{z}$ in all irreps, for example, see (Li et al., 2023). But this choice for jump operators preserves (parts of) dynamical sector structure, so there will be a NESS in each sector. In this way we generally get an initial state dependent NESS. If there is weak symmetry $P$ that $\mathcal{L}(P\rho P)=P\rho P$ or non-trivial strong symmetry (McDonald and Clerk, 2022; Buča and Prosen, 2012; Albert and Jiang, 2014; Zhang et al., 2020; Li et al., 2024) $S$ that commutes with all $\tilde{L}$ and $H$ (i.e., $S$ is in commutant of algebra of $\langle\{L_{k}\},\{L^{\dagger}_{k}\},H\rangle$, $[S,L_{k}]=[S,L^{\dagger}_{k}]=[S,H]=0$ for all $k$), we will have multiple NESS, because the Liouvillian is reducible (Evans, 1977; Frigerio, 1977; Baumgartner and Narnhofer, 2008; Zhang and Barthel, 2024) to independent sectors determined by invariant space of $S$.
@@
which is satisfied by $[\tilde{H},\tilde{L}_{k}]=\frac{\log(\lambda_{k})}{\beta}\tilde{L}_{k}$. So $\tilde{L}$s are ladder operators for $\tilde{H}$. We can thus take $\tilde{H}$ to be conserved quantities of the original Hamiltonian. Denote $\varepsilon=\frac{\log(\lambda_{k})}{\beta}$ and we have a tower of states of $\tilde{H}$ with energy $\{E_{0},E_{0}+\varepsilon,E_{0}+2\varepsilon,...\}$ and corresponding states are {|$\{\ket{\phi},\tilde{L}_{k}\ket{\phi},\tilde{L}_{k}^{2}\ket{\phi},...\}$ for $\tilde{H}\ket{\phi}=E_{0}\ket{\phi}$ and etc.. To define "temperature" we may let $\beta=\log(\lambda_{k})$ so $\varepsilon=1$. The $\tilde{H}$ behaves as Hamming distance from the root state |$\ket{\phi}$. Since all states in the same sector $\psi(\ket{\phi})$ has identical weight in NESS (5), there must be some additional constraints on jump operators. We can think of a dynamical sector as one equivalent class state $|\psi(\phi))$ and jump operators act as transitions between these states, $\tilde{L}|\psi(\phi))\equiv\tilde{L}\rho_{\psi(\phi)}\tilde{L}^{\dagger}$ and $(\psi(\phi)|\psi(\omega))=\text{Tr}(\rho_{\psi(\phi)}^{\dagger}\rho_{\psi(\omega)})$. Viewing dynamical sectors as vertices of a graph, then each edge corresponds to an application of jump operator $\tilde{L}$ and carries a positive weight $e^{-\beta}$ according to proportion of dissipation strength of $\tilde{L}$ and $\tilde{L}^{\dagger}$. The weight of a path is then defined as the product of all weights on the path and is related to energy of $\tilde{H}$. A key point in determining all $c_{\psi}$, or validity of the ansatz (5), relies on the fact that the graph being a wDAG generated by a set of $\{\tilde{L}_{i}\}$, and all path with the same source and end must have the same weight, which implies jump operators that carries different weights $\beta_{a}$, $\beta_{b}$ cannot generate the same path $(\psi(\omega)|\tilde{L}_{i}^{a\dagger}e^{iH\epsilon}\tilde{L}_{j}^{b}|\psi(\omega))=(\psi(\omega)|\tilde{L}_{i}^{a}e^{iH\epsilon}\tilde{L}_{j}^{b}|\psi(\omega))=0$ for arbitrary $\epsilon$, $\psi(\omega),i,j$ if $a\neq b$.
@@
Equilibrium environment.— To gain an intuitive view, let us consider the canonical XX model
@@
in contact of equilibrium environment with the jump operators given by pairs of $\sigma^{\pm}$ acting on the i-th site $L_{\text{Dis}}=\{L_{i,+}\equiv\sqrt{\gamma_{1}}\sigma_{i}^{+},\,L_{i,-}\equiv\sqrt{\gamma_{2}}\sigma_{i}^{-}\}$. Since they corresponds to raising and lowering of total magnetization, we know $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$ with $\beta=\log(\frac{\gamma_{1}}{\gamma_{2}})$ and the effective Hamiltonian is just the total magnetization $\tilde{H}=-\sum_{i}\sigma^{z}_{i}\equiv-\sigma^{z}_{\text{tot}}$ (Žnidarič, 2011).
@@
Non-equilibrium environment.— Next we study Fredkin spin chain (Salberger and Korepin, 2017) with boundary driving. The Fredkin spin chain is composed of local three-site spin interactions and described by
@@
We explicitly consider the Lindblad master equation of form (1) for system of length $N$ with jump operators $\text{Dis}=\{\sqrt{\gamma_{1}}L_{1},\sqrt{\gamma_{2}}L_{1}^{\dagger},\sqrt{\gamma_{3}}L_{N},\sqrt{\gamma_{4}}L_{N}^{\dagger}\}$ where $L_{i}=\sigma^{+}_{i}$ and we define $\beta_{l}\equiv\log(\frac{\gamma_{1}}{\gamma_{2}})$, $\beta_{r}\equiv\log(\frac{\gamma_{3}}{\gamma_{4}})$. Here we focus on the case with $\beta_{l}=-\beta_{r}=\beta$. Discussions on a general case is given in SM app . Without loss of generality we set $\gamma_{1}\geq\gamma_{2}$. Under this assumption particles flow into system from the left boundary and leave from the right driven by boundary jump terms.
@@
By "thermal equilibrium" we mean that the system is in equilibrium with bath, i.e., system exchanges particles or quantas of energy with bath according to DBC and NESS is a Gibbs state. By equilibrium statistical physics the "temperature" of system is naturally the same with bath. In terms of energy transport we can define temperature of bath $e^{\frac{1}{T_{\text{bath}}}}=\frac{\gamma_{1}}{\gamma_{2}}$ (Derrida, 2007), where we choose the energy unit to be 1. However in boundary driven system the temperature is generally distinct at different boundary. The fact that Fredkin spin chain can be in thermal equilibrium under DBC with two baths at different temperatures is surprising. We attribute this phenomenon to constraint of the dynamics originated from the model, which makes coupling of different conserved charges with different baths possible.
@@
Conclusions and outlook.— We unveiled that, by properly choosing jump operators, "finite temperature" Gibbs ensemble can emerge as the NESS for open quantum system described by Lindblad master equation. This construction is quite general for system with U(1) conservation laws. We explicitly construct NESS for models that possess one and two conserved charges, and find NESS can reach equilibrium even coupled with multiple bathes. An important future question is, how to construct similar Gibbs ensemble for system with non-Abelian symmetries like $\mathfrak{su}(2)$ of Heisenberg model or Temperley-Lieb model, a variant of pair-flip model with SU(3) symmetry. The difficulty lies in the construction of mappings between different irrep space. Besides, it is interesting to study time evolution under our construction, and examine to what extent quantum effect influence the dynamics, and how different placement or number of dissipation affect Liouvillian gap and the relaxation to the same NESS (Žnidarič, 2015; Han et al., 2024). We will leave these interesting questions to future work.

### v2

# Emergence of Gibbs ensembles as steady states in Lindbladian dynamics

@@ Abstract
We explicitly construct unique non-equilibrium steady state (NESS) of Lindblad master equation characterized by a Gibbs ensemble $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$, where the effective Hamiltonian $\tilde{H}$ consists only of $U(1)$ conserved charges of the original Hamiltonian. Specifically, when the original Hamiltonian has multiple charges, it is possible to couple them with bathes at different temperature respectively, but still leads to an equilibrium state. Multiple steady states arise if the number of bathes is less than the number of charges. To access the Gibbs NESS, the jump operators need to be properly chosen to fulfill quantum detailed balance condition (qDBC). These jump operators are ladder operators for $\tilde{H}$ and jump process they generate form a vertex-weighted directed acyclic graph (wDAG). By studying the XX model and Fredkin model, we showcase how Gibbs states emerge as steady states.
I Introduction
The Lindblad master equation is a paradigm for describing Markovian open quantum systems [1, 2, 3]. It represents the dynamics as an average of different trajectories, which consist of non-Hermitian evolution and stochastic quantum jumps, also known as "unraveling." In master equation form, the dynamics generator is
@@ I Introduction
where $L_{k}$s are jump operators and $H_{\text{eff}}=H-\frac{i}{2}\sum_{k}L_{k}^{\dagger}L_{k}$. The detailed balance condition (DBC) is not required in Lindblad dynamics, making it powerful and capable of characterizing non-equilibrium phenomena [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Recently with development of quantum computation, people regain interests in crafting quantum Gibbs samplers and preparing specific state utilizing Lindblad dynamics [18, 19, 20, 21]. In this work we make direct use of qDBC to derive exact Gibbs state that only depends on $U(1)$ conserved charges of original Hamiltonian.
DBC originates from classical statistical physics and guarantees an equilibrium stationary state. In a classical Markov process generated by W, the probability distribution of configuration set {c} satisfies $\frac{\mathrm{d}P(c)}{\mathrm{d}t}=\sum_{c^{\prime}\neq c}W(c^{\prime}\rightarrow c)P(c^{\prime})-W(c\rightarrow c^{\prime})P(c)$. For a stationary state satisfying classical DBC, $W(c^{\prime}\rightarrow c)P(c^{\prime})=W(c\rightarrow c^{\prime})P(c)$. DBC can also be interpreted as time-reversal symmetry, as it constrains the relative probabilities of configurations connected by a Markov process. In other words, starting from an arbitrary configuration, one can determine the probability of any other configuration without frustration. Besides the above intuitive condition, Kubo-Martin-Schwinger (KMS) condition [22] $\langle\hat{A}(t)\hat{B}(0)\rangle_{\text{eq}}=\langle\hat{B}(-t-i\beta)\hat{A}(0)\rangle_{\text{eq}}$ also acts as definition for thermal equilibrium state. It suggests a generalized "time evolution" acting on $\hat{B}$. qDBC extends this idea by requiring that this "time evolution" is also a completely positive and trace-preserving (CPTP) map, or a quantum Markov semigroup (QMS). Similar idea is applied to study properties of thermal states [23] and mixing time [24], to solve some models exactly [25, 26, 27] and a local thermal state is constructed [28]. In the context of Lindblad dynamics $\mathcal{L}(H,\{L_{k}\})$, the qDBC condition for an invertible (faithful) density matrix $\rho$ manifests as [29, 30]
@@
where $\tilde{L}_{k}=UL_{k}$, $U$ is an unitary that operates on space spanned by {$L_{k}$}. Liouvillian of this form that hosts a Gibbs NESS is often denoted as Davies generator [31].
@@
When jump operators are Hermitian, the identity matrix is always a NESS due to hermiticity of the Liouvillian. However, when jump operators are non-Hermitian and the Liouvillian is not diagonal-preserving, the NESS is typically not a diagonal equilibrium state and usually has finite stationary current, especially in the case of boundary driving [32]. Nevertheless, in this work we demonstrate that a diagonal equilibrium state can be constructed as a NESS in various models. Such a state satisfies qDBC, making it an equilibrium state that depends only on the conserved quantities of the original Hamiltonian.
@@
Our method introduced here relies on the coherent Hamiltonian having conservation laws, such that its Hilbert space is fractured into some disconnected pieces, which we denote as Krylov sectors or dynamical sectors. Besides conventional $U(1)$ conversational laws like particle conservation [33], it is also applicable to non-Abelian cases [34], dynamical constrained systems [35, 36] and systems with Hilbert space fragmentation (HSF) [37, 38, 39, 40].
II General theory
@@
Throughout this work we only consider separable Hilbert space with finite local dimension, such as spin system or fermion system. Our goal is to achieve a hierarchical structure of NESS that resembles equilibrium Gibbs distribution. Since a density matrix $\rho$ is positive and Hermitian, it can be expressed as $\rho=e^{-\beta\tilde{H}}$, where $\tilde{H}$ is a Hermitian matrix and $\beta\in\mathbb{R}$. To clarify Gibbs state, we construct $\tilde{H}$ by conserved quantities in original Hamiltonian and relate $\beta$ to temperature-alike quantity as in classical statistical physics. To realize this hierarchy, we pack states by taking conserved quantities as our packing principle. Since conserved quantities separate Hilbert space into disconnected sectors, every state sharing the same quantum numbers should have the same weight in $\rho$. An effective tool to characterize these conservation laws is commutant algebra. The commutant algebra formalism, as discussed in [38], is particularly useful for analyzing HSF, but in general it is not restricted to HSF cases. Commutant algebra formalism regards every invariant subspace of Hamiltonian as representation space for each irreducible representation (irrep) of bond algebra $\mathcal{A}$ generated by Hamiltonian terms $\{h_{\text{loc}}\}$ and identity operator $I$. The commutant algebra $\mathcal{C}$ consists of all operators $\{O\}$ of Hilbert space $\mathcal{H}$ that commute with all generators of $\mathcal{A}$, i.e., $[h_{\text{loc},i},O]=0$ for all $i$. The total Hilbert space can be decomposed by irreps of bond algebra and commutant algebra marked by $\lambda$, i.e., $\mathcal{H}=\bigoplus_{\lambda}\mathcal{H}^{\mathcal{A}}_{\lambda}\otimes\mathcal{H}^{\mathcal{C}}_{\lambda}$. If $\mathcal{H}^{\mathcal{C}}_{\lambda}$ has dimension $d$ larger than 1, then the Hamiltonian has degenerated sectors $H_{\lambda,1,...d}$. An operator in $\mathcal{A}$ has form $\oplus_{\lambda}h_{\lambda}\otimes\mathcal{I}$, and an operator in $\mathcal{C}$ has form $\oplus_{\lambda}\mathcal{I}\otimes c_{\lambda}$ where $h_{\lambda}$ and $c_{\lambda}$ is supported on $\mathcal{H}^{\mathcal{A}}_{\lambda}$ and $\mathcal{H}^{\mathcal{C}}_{\lambda}$, respectively.
@@
Now returning to qDBC (2), natural questions arise: how is $\tilde{H}$ related to conserved quantities of the original Hamiltonian and how can we find specific jump operators that satisfies qDBC? We will see they are related. First let us consider choice of jump operators. If $\lambda_{k}=1$, then a direct choice for $\tilde{L}_{k}$ to satisfy qDBC will be an element of bond algebra or a diagonal operator like $S^{z}$ in all irreps, for example, see [39]. But this choice for jump operators preserves (parts of) dynamical sector structure, so there will be a NESS in each sector. In this way we generally get an initial state dependent NESS. If there is weak symmetry $P$ that $\mathcal{L}(P\rho P)=P\rho P$ or non-trivial strong symmetry [41, 33, 42, 34, 43] $S$ that commutes with all $\tilde{L}$ and $H$ (i.e., $S$ is in commutant of algebra of $\langle\{L_{k}\},\{L^{\dagger}_{k}\},H\rangle$, $[S,L_{k}]=[S,L^{\dagger}_{k}]=[S,H]=0$ for all $k$), we will have multiple NESS, because the Liouvillian is reducible [44, 45, 46, 47] to independent sectors determined by invariant space of $S$.
@@
which is satisfied by $[\tilde{H},\tilde{L}_{k}]=\frac{\log(\lambda_{k})}{\beta}\tilde{L}_{k}$. So $\tilde{L}$s are ladder operators for $\tilde{H}$. We can thus take $\tilde{H}$ to be conserved quantities of the original Hamiltonian. Denote $\varepsilon=\frac{\log(\lambda_{k})}{\beta}$ and we have a tower of states of $\tilde{H}$ with energy $\{E_{0},E_{0}+\varepsilon,E_{0}+2\varepsilon,...\}$ and corresponding states are {|$\{\ket{\phi},\tilde{L}_{k}\ket{\phi},\tilde{L}_{k}^{2}\ket{\phi},...\}$ for $\tilde{H}\ket{\phi}=E_{0}\ket{\phi}$ and etc.. To define "temperature" we may let $\beta=\log(\lambda_{k})$ so $\varepsilon=1$. The $\tilde{H}$ behaves as Hamming distance from the root state |$\ket{\phi}$. Since all states in the same sector $\psi(\ket{\phi})$ has identical weight in NESS (5), there must be some additional constraints on jump operators. We can think of a dynamical sector as one equivalent class state $|\psi(\phi))$ and jump operators act as transitions between these states, $\tilde{L}|\psi(\phi))\equiv\tilde{L}\rho_{\psi(\phi)}\tilde{L}^{\dagger}$ and $(\psi(\phi)|\psi(\omega))=\text{Tr}(\rho_{\psi(\phi)}^{\dagger}\rho_{\psi(\omega)})$. Viewing dynamical sectors as vertices of a graph, then each edge corresponds to an application of jump operator $\tilde{L}$ and carries a positive weight $e^{-\beta}$ according to proportion of dissipation strength of $\tilde{L}$ and $\tilde{L}^{\dagger}$. The weight of a path is then defined as the product of all weights on the path and is related to energy of $\tilde{H}$. A key point in determining all $c_{\psi}$, or validity of the ansatz (5), relies on the fact that the graph being a weighted directed acyclic graph (wDAG) generated by a set of $\{\tilde{L}_{i}\}$, and all path with the same source and end must have the same weight, which implies jump operators that carries different weights $\beta_{a}$, $\beta_{b}$ cannot generate the same path $(\psi(\omega)|\tilde{L}_{i}^{a\dagger}e^{iH\epsilon}\tilde{L}_{j}^{b}|\psi(\omega))=(\psi(\omega)|\tilde{L}_{i}^{a}e^{iH\epsilon}\tilde{L}_{j}^{b}|\psi(\omega))=0$ for arbitrary $\epsilon$, $\psi(\omega),i,j$ if $a\neq b$.
@@
III Equilibrium environment
To gain an intuitive view, let us consider the canonical XX model
@@ III Equilibrium environment
in contact of equilibrium environment with the jump operators given by pairs of $\sigma^{\pm}$ acting on the i-th site $L_{\text{Dis}}=\{L_{i,+}\equiv\sqrt{\gamma_{1}}\sigma_{i}^{+},\,L_{i,-}\equiv\sqrt{\gamma_{2}}\sigma_{i}^{-}\}$. Since they corresponds to raising and lowering of total magnetization, we know $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$ with $\beta=\log(\frac{\gamma_{1}}{\gamma_{2}})$ and the effective Hamiltonian is just the total magnetization $\tilde{H}=-\sum_{i}\sigma^{z}_{i}\equiv-\sigma^{z}_{\text{tot}}$ [48].
@@
IV Non-equilibrium environment
Next we study Fredkin spin chain [49] with boundary driving. The Fredkin spin chain is composed of local three-site spin interactions and described by
@@ IV Non-equilibrium environment
We explicitly consider the Lindblad master equation of form (1) for system of length $N$ with jump operators $\text{Dis}=\{\sqrt{\gamma_{1}}L_{1},\sqrt{\gamma_{2}}L_{1}^{\dagger},\sqrt{\gamma_{3}}L_{N},\sqrt{\gamma_{4}}L_{N}^{\dagger}\}$ where $L_{i}=\sigma^{+}_{i}$ and we define $\beta_{l}\equiv\log(\frac{\gamma_{1}}{\gamma_{2}})$, $\beta_{r}\equiv\log(\frac{\gamma_{3}}{\gamma_{4}})$. Here we focus on the case with $\beta_{l}=-\beta_{r}=\beta$. Discussions on a general case is given in SM [50]. Without loss of generality we set $\gamma_{1}\geq\gamma_{2}$. Under this assumption particles flow into system from the left boundary and leave from the right driven by boundary jump terms.
@@
By "thermal equilibrium" we mean that the system is in equilibrium with bath, i.e., system exchanges particles or quantas of energy with bath according to DBC and NESS is a Gibbs state. By equilibrium statistical physics the "temperature" of system is naturally the same with bath. In terms of energy transport we can define temperature of bath $e^{\frac{1}{T_{\text{bath}}}}=\frac{\gamma_{1}}{\gamma_{2}}$ [5], where we choose the energy unit to be 1. However in boundary driven system the temperature is generally distinct at different boundary. The fact that Fredkin spin chain can be in thermal equilibrium under DBC with two baths at different temperatures is surprising. Actually it is general for systems with non-Abelian symmetry or multiple U(1) conserved charges, which makes coupling of different conserved charges with different baths possible. In this way it is also straight forward to have multiple steady states if the number of bathes is less than the number of charges. For example, in Fig.1 deleting all red(blue) arrows will generate 3 independent Gibbs states.
@@
V Conclusions and outlook
We unveiled that, by properly choosing jump operators, "finite temperature" Gibbs ensemble can emerge as the NESS for open quantum system described by Lindblad master equation. This construction is quite general for system with U(1) conservation laws. We explicitly construct NESS for models that possess one and two conserved charges, and find NESS can reach equilibrium even coupled with multiple bathes. An important future question is, how to construct similar Gibbs ensemble for system with non-Abelian symmetries like $\mathfrak{su}(2)$ of Heisenberg model or Temperley-Lieb model, a variant of pair-flip model with SU(3) symmetry. The difficulty lies in the construction of mappings between different irrep space. Besides, it is interesting to study time evolution under our construction, and examine to what extent quantum effect influence the dynamics, and how different placement or number of dissipation affect Liouvillian gap and the relaxation to the same NESS [51, 52]. Moreover, since we utilize Lindblad master equation which describes Markovian dynamics, it is also interesting how to apply our methods in non-Markovian cases, where there are some pioneer work [53, 54]. We will leave these questions to future work.